The size of the screen of a television set is usually given in terms of its diagonal . Find the length of the diagonal of the screen 15 cm wide and 20 cm long.
Answer by amoresroy(361)

The size of the screen of a television set is usually given in terms of its diagonal . Find the length of the diagonal of the screen 15 cm wide and 20 cm long.
Let d = the length of the diagonal of the screen in cm
by phytagorean formula
d^2 = 15^2 + 20^2
d^2 = 225 + 400
d = (625)^1/2
d = 25
The length of the diagonal of the screen is 25 cm.